OpenStudy (anonymous):

Find three consecutive integers such that the sum of the second and the third is 17.

OpenStudy (anonymous):

7,8,9

OpenStudy (anonymous):

the options are a. x + (x + 1) + (x + 2) = 17 b. (x + 1) + (x + 2) = 17 c. x + (x + 2) + (x + 4) = 17 d. 2(x + 2) + 3(x + 4) = 17

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks for the help! :)